OpenStudy (anonymous):

The parabolic arch of a bridge over a one-way road can be described by the equation y = -x^2 + 6, where x and y ar emeasured in metres. A transpor truck is 3m wide and 3.5m high. What is the maximum clearance of the truck's top corners if it passes under the bridge?

OpenStudy (anonymous):

This is an Optimization problem.

OpenStudy (dumbcow):

so you want the max distance from the top corner of truck to arch ??

OpenStudy (dumbcow):

distance from top center of arch to top corner of truck is 2.915 m

OpenStudy (anonymous):

Yes How were you able to work that out?

OpenStudy (dumbcow):

d^2 = 2.5^2 + 1.5^2

OpenStudy (anonymous):

How did you come to this equation?

OpenStudy (dumbcow):

truck is 3.5m so vertical distance from top is (6-3.5) = 2.5 truck is 3m wide, 1.5m on either side of center line imagine a triangle formed from right side of top of truck to top center of arch d is hypotenuse and is the longest distance from arch to truck top corner